Solve. 5/2x+1/4x=11/4+x
Answer by oberobic(2304)

5/2x + 1/4x = 11/4 + x
.
These fractions are painful. But we can treat them all as quarters.
.
Recall
5/2 = 10/4
Also,
x = 1/1x = 2/2x = 3/3x = 4/4x...
.
10/4x + 1/4x = 11/4 + 4/4x
.
11/4x = 11/4 + 4/4x
.
Subtract 4/4x from both sides
7/4x = 11/4
.
Multiply both sides by 4
7x = 11
.
Divide both sides by 7
x = 11/7
.